# Non-isogenous superelliptic jacobians II

### v1

@@ 1. Definitions, notations, statements
This paper is a follow up of [18, 21] and we use its (more or less standard) notation. (See also [16, 20, 21].) In particular, $\ell$ is an odd prime, ${\mathbb{F}}_{\ell}$ the corresponding (finite) prime field of characteristic $\ell$. We write ${\mathbb{Z}}_{\ell}$ and ${\mathbb{Q}}_{\ell}$ for the ring of $\ell$-adic integers and the field ${\mathbb{Q}}_{\ell}$ of $\ell$-adic numbers respectively. Let us fix a primitive $\ell$th root of unity
@@
${\mathbb{Z}}[\zeta_{q}]=\sum_{i=0}^{\phi(q)-1}{\mathbb{Z}}\cdot\zeta_{q}^{i}$
@@
Let $K$ be a field with $\mathrm{char}(K)\neq\ell$. Let us fix an algebraic closure $\bar{K}$ of $K$ and write $\mathrm{Gal}(K)=\mathrm{Aut}(\bar{K}/K)$ for its group of its $K$-linear automorpisms. In what follows we always assume that $K$ contains a primitive $\ell$th root of unity, say, $\zeta$.
Let $f(x)\in K[x]$ be a polynomial with coefficients in $K$, of degree $n$ not diivisible by $\ell$, and without repeated roots. We write $\mathfrak{R}_{f}\subset\bar{K}$ for the $n$-element set of roots of $f(x)$, $K(\mathfrak{R}_{f})$ for the splitting field of $f(x)$ and $\mathrm{Gal}(f/K)$ for the Galois group
@@
We write $C_{f,\ell}$ for the smooth projective model of the plane affine curve $y^{\ell}=f(x)$. It is well known [10, 11] that the genus $g(C_{f,\ell})$ of $C_{f,\ell}$ is $(\ell-1)(n-1)/2$ if $\ell$ does not divide $n$. Further we assume that $\ell\nmid n$. The map
@@
Let $J^{(f,\ell)}$ be the jacobian of $C_{f,\ell}$; it is a $(\ell-1)(n-1)/2$-dimensional abelian variety that is defined over $K$. We write $\mathrm{End}(J^{(f,\ell)})$ for the ring of $\bar{K}$-endomorphisms of $J^{(f,\ell)}$ and $\mathrm{End}^{0}((J^{(f,\ell)})=\mathrm{End}(J^{(f,\ell)})\otimes{\mathbb{Q}}$ for the corresponding endomorphism algebra of $J^{(f,\ell)}$. By functoriality, $\delta_{\ell}$ induces a $K$-automorphism of $J^{(f,\ell)}$, which we still denote by $\delta_{q}$. It is known ([10, p. 149], [11, p. 448]; see also [19]) that
@@
which is a ring embedding ([10, p. 149], [11, p. 448]; see also [19]). This implies that the subring ${\mathbb{Z}}[\delta_{\ell}]$ of $\mathrm{End}(J^{(f,\ell)})$ generated by $\delta_{\ell}$ is isomorphic to ${\mathbb{Z}}[\zeta_{\ell}]$. It follows that the ${\mathbb{Q}}$-subalgebra
@@
generated by $\delta_{\ell}$ is canonically isomorphic to the $\ell$th cyclotomic field ${\mathbb{Q}}(\zeta_{\ell})$ and therefore has ${\mathbb{Q}}$-dimension $\ell-1$.
Throughout this paper, $\ell$ is a prime, $n\geq 3$ is an integer not divisible by $\ell$, and $f(x)$ and $h(x)$ are degree $n$ polynomials with coefficients in $K$ and without repeated roots,
@@
are the corresponding genus $(n-1)(\ell-1)/2$ superelliptic curves over $K$, whose jacobians we denote by $J^{(f,\ell)}$ and $J^{(h,\ell)}$, respectively. These jacobians are $(n-1)(\ell-1)/2$-dimensional abelian varieties defined over $K$. Hereafter $\ell$ is an odd prime that does not divide $n$.
@@
Then both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $n$.
The next assertion may be viewed as a partial generalization of Theorem 1.1 to the case of arbitrary odd $n$.
@@
If the corresponding superelliptic jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then there is a prime divisor $r$ of $n$ such that both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $r$.
@@
If the conditions of Theorem 1.2 hold then $h(x)$ is reducible over $K$, see [21, Remark 1.4].
@@
By Eisenstein criterion, $f(x)$ is irreducible over $K$ (recall a prime $2$ is unramified in ${\mathbb{Q}}(\zeta_{\ell})=K$, because $\ell$ is odd) while $h(x)$ is obviously reducible over $K$. Let $\ell$ be an odd prime that does not divide $n$. The curves $C_{f,\ell}$ and $C_{h,\ell}$ are obviously isomorphic over $\bar{K}$. They both admit periodic $\bar{K}$-automorphisms $\tilde{\delta}_{n}$ of order $n$ defined by the (same) formula
@@
is a primitive $n$th root of unity. This implies that their jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are abelian varieties over $\bar{K}$ that are isomorphic over $\bar{K}$ and admit periodic automorphisms of (multiplicative) order $n$ that we continue to denote by $\tilde{\delta}_{n}$. It follows that if $r$ is any prime divisor of $n$ then both $J^{(f,\ell)}$ and $J^{(h,\ell)}$ admit automorphisms $\left(\tilde{\delta}_{n}\right)^{n/r}$ of multiplicative order $r$ that may be viewed as invertible elements of multiplicative order $r$ in $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$.
@@
Let $n\geq 5$ be an odd integer and $\ell$ an odd prime not dividing $n$. Let us put $K={\mathbb{Q}}(\zeta_{n})$.
@@
Then it follows from Theorem 1.2 that $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are not isogenous over $\bar{K}=\bar{{\mathbb{Q}}}$ (and therefore even over $\bar{{\mathbb{C}}}$). Since $J^{(f,\ell)}$ is absolutely simple and has the same dimension as $J^{(h,\ell)}$, it follows that every ${\mathbb{C}}$-homomorphism between $J^{(f,\ell)}$ and $J^{(h,\ell)}$ is zero.
@@
Recall that $n\geq 5$ and therefore $\mathrm{A}_{n}$ is a simple non-abelian group that coincides with the commutator subgroup of $\mathrm{S}_{n}$; in addition, $\mathrm{A}_{n}$ is a maximal subgroup of $\mathrm{S}_{n}$. Since $K={\mathbb{Q}}(\zeta_{\ell})$ is an abelian extension of ${\mathbb{Q}}$, the Galois groups $\mathrm{Gal}(f_{1}/K)$ and $\mathrm{Gal}(f_{2}/K)$ are either $\mathrm{S}_{n}$ or $\mathrm{A}_{n}$; in particular, both $f_{1}$ and $f_{2}$ remain irreducible over $K$.
@@
The paper is organized as follows. In Section 2 we recall basic facts about Galois properties of points of order $\ell$ on superelliptic jacobians. We also state Theorem 2.2 that is a slightly stronger version of Theorem 1.2. Section 3 contains the proof of Theorem 2.2. We prove Theorem 1.1 in Section 4.
@@ 2. Points of order $\ell$ on superelliptic jacobians
Let $K_{s}\subset\bar{K}$ be the separable algebraic closure of $K$. Let $X$ be a positive-dimensional abelian variety over $K$. If $d$ is a positive integer then we write $X[d]$ for the kernel of multiplication by $d$ in $X(\bar{K})$. Recall ([7, Sect. 6], [5, Sect. 8, Remark 8.4]) that if $d$ is not divisible by $\mathrm{char}(K)$ then $X[d]$ is a $\mathrm{Gal}(K)$-submodule of $X(K_{s})$; in addition, $X[d]$ is isomorphic as a commutative group to $({\mathbb{Z}}/d{\mathbb{Z}})^{2\mathrm{dim}(X)}$.
Let $K(X[d])$ be the field of definition of all torsion points of order dividing $d$ on $X$. It is well known [5, Remark 8.4] that $K(X[d])$ lies in $K_{s}$ and is a finite Galois extension of $K$. Let us put
$\tilde{G}_{d,X}:=\mathrm{Gal}(K(X[d])/K).$
One may view $\tilde{G}_{d,X}$ as a certain subgroup of $\mathrm{Aut}_{{\mathbb{Z}}/d{\mathbb{Z}}}(X[d])$ and $X[d]$ as a faithful $\tilde{G}_{d,X}$-module. In addition, the structure of the $\mathrm{Gal}(K)$-module on $X[d]$ is induced by the canonical (continuous) surjective group homomorphism
$\tilde{\rho}_{d,X}:\mathrm{Gal}(K)\twoheadrightarrow\mathrm{Gal}(K(X[d])/K)=\tilde{G}_{d,X}.$
For example, if $d$ is a prime $\ell\neq\mathrm{char}(K)$ then $X[\ell]$ is a $2\mathrm{dim}(X)$-dimensional vector space over the $\ell$-element prime field ${\mathbb{F}}_{\ell}={\mathbb{Z}}/\ell{\mathbb{Z}}$ and the inclusion $\tilde{G}_{\ell,X}\subset\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(X[\ell])$ defines a faithful linear representation of the group $\tilde{G}_{\ell,X}$ in the vector space $X[\ell]$ over ${\mathbb{F}}_{\ell}$.
@@
If $d\geq 3$ is an integer that is every ${\mathbb{C}}$-homomorphism between $J^{(f,\ell)}$ and $J^{(h,\ell)}$ is zero.not divisible by $\mathrm{char}(K)$ then all the endomorphisms of $X$ are defined over $K(X[d])$, by a theorem of A. Silverberg [14, Th. 2.4]. (See also [3, 12, 2].)
@@
Let $Y$ be a $(n-1)(\ell-1)/2$-dimensional abelian variety over $K$ such that the order of $\tilde{G}_{\ell,Y}$ is prime to $n$. (E.g., $K(Y[\ell])=K$ or this order is a power of $\ell$.)
@@
Then there is an odd prime $r$ dividing $n$ such that both endomorphism algebras $\mathrm{End}^{0}(X)$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$.

### v4

@@ 1. Definitions, notations, statements
This paper is a follow up of [20, 23] and we use their (more or less standard) notation. (See also [18, 22, 23].) In particular, $\ell$ is an odd prime, ${\mathbb{F}}_{\ell}$ the corresponding (finite) prime field of characteristic $\ell$. We write ${\mathbb{Z}}_{\ell}$ and ${\mathbb{Q}}_{\ell}$ for the ring of $\ell$-adic integers and the field ${\mathbb{Q}}_{\ell}$ of $\ell$-adic numbers respectively. Let us fix a primitive $\ell$th root of unity
@@
${\mathbb{Z}}[\zeta_{\ell}]=\sum_{i=0}^{\phi(q)-1}{\mathbb{Z}}\cdot\zeta_{\ell}^{i}$
@@
Let $K$ be a field with $\mathrm{char}(K)\neq\ell$. Let us fix an algebraic closure $\bar{K}$ of $K$ and write $\mathrm{Gal}(K)=\mathrm{Aut}(\bar{K}/K)$ for the group of its $K$-linear automorpisms. In what follows we always assume that $K$ contains a primitive $\ell$th root of unity, say, $\zeta$. Let $K_{s}\subset\bar{K}$ be the separable algebraic closure of $K$. The subfield $K_{s}$ of $\bar{K}$ is the natural group homomorphism (restriction map)
$\mathrm{Gal}(K)=\mathrm{Aut}(\bar{K}/K)\to\mathrm{Aut}(K_{s}/K)=\mathrm{Gal}(K_{s}/K)$
is a group isomorphism.
Let $X$ and $Y$ be abelian varieties over $K$. By a theorem of Chow [2, Th. 3.19]), all $\bar{K}$-endomorphisms between $X$ and $Y$ are defined over $K_{s}$. In particular, all endomorphisms of $X$ are defined over $K_{s}$. A more precise information about the field of definition of all endomorphisms of $X$ is given by a theorem of A. Silverberg [16] (see Remark 2.1 below).
Let $n\geq 3$ be an odd integer. Throughout the paper, we assume that $\ell$ does not divide $n$.
Let $f(x)\in K[x]$ be a polynomial with coefficients in $K$, of degree $n$ and without repeated roots. We write $\mathfrak{R}_{f}\subset\bar{K}$ for the $n$-element set of roots of $f(x)$, $K(\mathfrak{R}_{f})$ for the splitting field of $f(x)$ and $\mathrm{Gal}(f/K)$ for the Galois group
@@
We write $C_{f,\ell}$ for the smooth projective model of the plane affine curve $y^{\ell}=f(x)$. It is well known [12, 13] that the genus $g(C_{f,\ell})$ of $C_{f,\ell}$ is $(\ell-1)(n-1)/2$. The map
@@
Let $J^{(f,\ell)}$ be the jacobian of $C_{f,\ell}$; it is a $(\ell-1)(n-1)/2$-dimensional abelian variety that is defined over $K$. We write $\mathrm{End}(J^{(f,\ell)})$ for the ring of $\bar{K}$-endomorphisms of $J^{(f,\ell)}$ and $\mathrm{End}^{0}((J^{(f,\ell)})=\mathrm{End}(J^{(f,\ell)})\otimes{\mathbb{Q}}$ for the corresponding endomorphism algebra of $J^{(f,\ell)}$. By functoriality, $\delta_{\ell}$ induces a $K$-automorphism of $J^{(f,\ell)}$, which we still denote by $\delta_{\ell}$. It is known ([12, p. 149], [13, p. 448]; see also [21]) that
@@
which is a ring embedding ([12, p. 149], [13, p. 448]; see also [21]). (Here $1\in{\mathbb{Z}}[\zeta_{\ell}]$ goes to the identity automorphism $\mathrm{1}_{J^{(f,\ell)}}$ of $J^{(f,\ell)}$.) This implies that the subring ${\mathbb{Z}}[\delta_{\ell}]$ of $\mathrm{End}(J^{(f,\ell)})$ generated by $\delta_{\ell}$ is isomorphic to ${\mathbb{Z}}[\zeta_{\ell}]$. It follows that the ${\mathbb{Q}}$-subalgebra
@@
generated by $\delta_{\ell}$ is canonically isomorphic to the $\ell$th cyclotomic field ${\mathbb{Q}}(\zeta_{\ell})$ and therefore
$\mathrm{dim}_{{\mathbb{Q}}}({\mathbb{Q}}[\delta_{\ell}])=\ell-1.$
Let $f(x)$ and $h(x)$ be degree $n$ polynomials with coefficients in $K$ and without repeated roots. Let
@@
be the corresponding genus $(n-1)(\ell-1)/2$ superelliptic curves over $K$, whose jacobians we denote by $J^{(f,\ell)}$ and $J^{(h,\ell)}$, respectively. These jacobians are $(n-1)(\ell-1)/2$-dimensional abelian varieties defined over $K$. Assuming that the Galois properties of roots of $f(x)$ and $h(x)$ are distinct (see below), we will prove that if the abelian varieties $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then they admit an “additional symmetry”, i.e., the inclusion (3) is not an equality.
@@
If the corresponding superelliptic jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $n$. In addition,
$\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(f,\ell)})\right)=\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(h,\ell)})\right)\geq$
$(\ell-1)(n-1)=2\ \mathrm{dim}(J^{(f,\ell)})=2\ \mathrm{dim}(J^{(h,\ell)}).$
The next assertion may be viewed as a partial generalization of Theorem 1.1 to the case of an arbitrary odd $n\geq 3$.
@@
If the corresponding superelliptic jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then there is a prime divisor $r$ of $n$ such that both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $r$. In addition,
$\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(f,\ell)})\right)=\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(h,\ell)})\right)\geq(\ell-1)(r-1).$
@@
If the conditions of Theorem 1.2 hold then $h(x)$ is reducible over $K$, see [23, Remark 1.4].
Corollary 1.4.
Suppose that $n$ and $\ell$ are distinct odd primes. Let $f(x),h(x)\in K[x]$ be degree $n$ polynomials without repeated roots. Suppose that $f(x)$ is irreducible over $K$ and $\mathrm{Gal}(\mathfrak{R}_{f})$ is a doubly reansitive permutation group of $\mathfrak{R}_{f}$. Assume also that $\mathrm{Gal}(h/K)$ is a cyclic group of order $n$. If the corresponding superelliptic jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $n$. In addition,
$\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(f,\ell)})\right)=\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(h,\ell)})\right)\geq$
$(\ell-1)(n-1)=2\ \mathrm{dim}(J^{(f,\ell)})=2\ \mathrm{dim}(J^{(h,\ell)}).$
@@
By the $2$-dic Eisenstein criterion, $f(x)$ is irreducible over $K$ (recall that prime $2$ is unramified in ${\mathbb{Q}}(\zeta_{\ell})=K$, because $\ell$ is odd) while $h(x)$ is obviously reducible over $K$. Let $\ell$ be an odd prime that does not divide $n$. The curves $C_{f,\ell}$ and $C_{h,\ell}$ are obviously isomorphic over $\bar{K}$. They both admit periodic $\bar{K}$-automorphisms $\tilde{\delta}_{n}$ of order $n$ defined by the (same) formula
@@
is a primitive $n$th root of unity. This implies that their jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are abelian varieties over $\bar{K}$ that are isomorphic over $\bar{K}$ and admit periodic automorphisms of order $n$ that we continue to denote by $\tilde{\delta}_{n}$. It follows that if $r$ is any prime divisor of $n$ then both $J^{(f,\ell)}$ and $J^{(h,\ell)}$ admit automorphisms $\left(\tilde{\delta}_{n}\right)^{n/r}$ of multiplicative order $r$ that may be viewed as invertible elements of multiplicative order $r$ in $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$.
@@
Let $n\geq 5$ be an odd integer and $\ell$ an odd prime not dividing $n$. Let us put $K={\mathbb{Q}}(\zeta_{\ell})$.
@@
Then it follows from Theorem 1.2 that $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are not isogenous over $\bar{K}=\bar{{\mathbb{Q}}}$ (and therefore even over $\bar{{\mathbb{C}}}$, in light of a theorem of Chow [2, Th. 3.19]). Since $J^{(f,\ell)}$ is absolutely simple and has the same dimension as $J^{(h,\ell)}$, it follows that every ${\mathbb{C}}$-homomorphism between $J^{(f,\ell)}$ and $J^{(h,\ell)}$ is zero.
@@
Recall that $n\geq 5$ and therefore $\mathrm{A}_{n}$ is a simple non-abelian group that coincides with the commutator subgroup of $\mathrm{S}_{n}$; in addition, $\mathrm{A}_{n}$ is a maximal subgroup of $\mathrm{S}_{n}$. Since $K:={\mathbb{Q}}(\zeta_{\ell})$ is an abelian extension of ${\mathbb{Q}}$, the Galois group $\mathrm{Gal}(f_{k}/K)$ is either $\mathrm{S}_{n}$ or $\mathrm{A}_{n}$ (for $k=1,2$). In particular, both $f_{1}$ and $f_{2}$ remain irreducible over $K$.
@@
Remark 1.8.
Examples 1.6 and 1.7 illustrate the title of this paper, whose aim is to provide a criterion for certain superelliptic jacobians not to be isogenous.
Remark 1.9.
In light of Theorem of Chow cited above, the assertions of Theorem 1.1, 1.2 (and of Theorem 2.2 below), and Corollary 1.4 remain true if one replaces $\bar{K}$ by $K_{s}$.
The paper is organized as follows. In Section 2 we recall basic facts about Galois properties of points of order $\ell$ on superelliptic jacobians. We also state Theorem 2.2 that is a slightly stronger version of Theorem 1.2. Section 3 contains the proof of Theorem 2.2. We prove Theorem 1.1 in Section 4. Corollary 1.4 is proven in Section 5.
Acknowledgments. I am deeply grateful to the referee for helpful comments.
@@ 2. Points of order $\ell$ on superelliptic jacobians
Recall that $K_{s}\subset\bar{K}$ is the separable algebraic closure of $K$. Let $X$ be a positive-dimensional abelian variety over $K$. We write $\mathrm{End}(X)$ for the ring of all $\bar{K}$-endomorphisms of $X$ and $\mathrm{End}^{0}(X):=\mathrm{End}(X)\otimes{\mathbb{Q}}$ for the corresponding endomorphism algebra of $X$, which is a semisimple finite-dimensional ${\mathbb{Q}}$-algebra. We write $\mathrm{1}_{X}$ for the identity automorphism of $X$ that is may be viewed as the identity element of the ${\mathbb{Q}}$-algebra $\mathrm{End}^{0}(X)$.
If $d$ is a positive integer then we write $X[d]$ for the kernel of multiplication by $d$ in $X(\bar{K})$. Recall ([9, Sect. 6], [7, Sect. 8, Remark 8.4]) that if $d$ is not divisible by $\mathrm{char}(K)$ then $X[d]$ is a $\mathrm{Gal}(K)$-submodule of $X(K_{s})$; in addition, $X[d]$ is isomorphic as a commutative group to $({\mathbb{Z}}/d{\mathbb{Z}})^{2\mathrm{dim}(X)}$.
Let $K(X[d])$ be the field of definition of all torsion points of order dividing $d$ on $X$. It is well known [7, Remark 8.4] that $K(X[d])$ lies in $K_{s}$ and is a finite Galois extension of $K$. Let us put
$\tilde{G}_{d,X,K}:=\mathrm{Gal}(K(X[d])/K).$
One may view $\tilde{G}_{d,X,K}$ as a certain subgroup of $\mathrm{Aut}_{{\mathbb{Z}}/d{\mathbb{Z}}}(X[d])$ and $X[d]$ as a faithful $\tilde{G}_{d,X,K}$-module. In addition, the structure of the $\mathrm{Gal}(K)$-module on $X[d]$ is induced by the canonical (continuous) surjective group homomorphism
$\tilde{\rho}_{d,X}:\mathrm{Gal}(K)\twoheadrightarrow\mathrm{Gal}(K(X[d])/K)=\tilde{G}_{d,X,K}.$
For example, if $d$ is a prime $\ell\neq\mathrm{char}(K)$ then $X[\ell]$ is a $2\mathrm{dim}(X)$-dimensional vector space over the field ${\mathbb{F}}_{\ell}={\mathbb{Z}}/\ell{\mathbb{Z}}$, and the inclusion $\tilde{G}_{\ell,X,K}\subset\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(X[\ell])$ defines a faithful linear representation of the group $\tilde{G}_{\ell,X,K}$ in the vector space $X[\ell]$ over ${\mathbb{F}}_{\ell}$.
@@
If $d\geq 3$ is an integer not divisible by $\mathrm{char}(K)$ then all the endomorphisms of $X$ are defined over $K(X[d])$, by a theorem of A. Silverberg [16, Th. 2.4]. (See also [4, 14, 3].)
@@
Let $Y$ be an abelian variety over $K$ such that the order of $\tilde{G}_{\ell,Y,K}$ is prime to $n$. (E.g., $K(Y[\ell])=K$ or this order is a power of $\ell$.)
@@
Then there is an odd prime $r$ dividing $n$ such that both endomorphism algebras $\mathrm{End}^{0}(X)$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$. In addition,
$\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(J^{(f,\ell)})\right)=\mathrm{dim}_{{\mathbb{Q}}}\left(\mathrm{End}^{0}(Y)\right)\geq(\ell-1)(r-1).$
Remark 2.3.
Recall that the order of $\tilde{G}_{\ell,Y,K}$ coincides with the degree $[K(Y[\ell]):K]$ of the Galois extension $K(Y[\ell])/K$.